OpenStudy (anonymous):

Can someone help me calculate this? your mass (kg) × acceleration of gravity (9.81m/s/s) = your weight (N) your weight (N) × height of board (m) = work done in stepping up onto board

OpenStudy (anonymous):

Your mass (68) = your mass in lbs 150/2.2 ) Use a bathroom scale to get you mass in lbs.

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

Would it be (68) x (9.81,/s/s) = 150 (150) x (24) =_______ Did I arrange that correctly?

OpenStudy (anonymous):

The height of the board is 24cm

OpenStudy (agent0smith):

24cm = 0.24m always use metres in calculations

OpenStudy (anonymous):

Ooops, yeah. So I did, set it up right though?

OpenStudy (agent0smith):

Looks like it

OpenStudy (agent0smith):

68 times 9.8 is definitely not 150. You should be able to tell, by estimating, noticing that 68*10 = 680... 150 is not close

OpenStudy (anonymous):

Nuu 150 is my weight, and 68 is my weight in kg

OpenStudy (anonymous):

8.9 is the time it took me so replace that with 9.8.

OpenStudy (agent0smith):

(68) x (9.81,/s/s) =???

OpenStudy (anonymous):

667.08

OpenStudy (agent0smith):

That's your weight in Newtons. Multiply it by the height

OpenStudy (anonymous):

16009.92?

OpenStudy (agent0smith):

No, scroll back to my very first post

OpenStudy (anonymous):

Wait, I don't see what I did wrong, I multiplied the height.

OpenStudy (agent0smith):

Again, see my first post...

OpenStudy (anonymous):

Ahh! 160.0992

OpenStudy (anonymous):

(68) x (8.9,/s/s) = 667.08 (667.08) x (0.24) = 160.0992

OpenStudy (agent0smith):

Gravity is 9.8 not 8.9

OpenStudy (anonymous):

Oh I thought it meant the acceleration of the speed. . .Between the reps I was doing. Eh, confused.

OpenStudy (anonymous):

Agent?

OpenStudy (agent0smith):

It looks correct, as long as you used 9.8 and not 8.9.

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (anonymous):

(68) x (8.9,/s/s) = 666.4 (666.4) x (0.24) =159.936

OpenStudy (agent0smith):

You are still using 8.9 for some reason...

OpenStudy (anonymous):

The time has to be different otherwise my trials will be the same.

OpenStudy (agent0smith):

Look what you originally posted: your mass (kg) × acceleration of gravity (9.81m/s/s) = your weight (N) your weight (N) × height of board (m) = work done in stepping up onto board

OpenStudy (agent0smith):

It's not time, it's gravity. You posted it right at the top.

OpenStudy (anonymous):

I dun know it wouldn't make sense then, because I have two trials, if I used the same number it would give me the same result, can't be right. I'm thinking it means my time and speed.

OpenStudy (anonymous):

Trial Height of board (m)24(M)>0.24 Weight (n)150(N) Work for 1 repetition (J)129.2(J) Time for 20 repetitions (S)8.9(s) Work for 20 repetitions (J)2584(J) Power for 20 repetitions (J/s)290.337(J/s)